OpenStudy (anonymous):

Solve the equation 3x = 3(x – 4) + 12

OpenStudy (ddcamp):

\[3x = 3(x-4) + 12 \rightarrow 3x = 3x - 12 + 12 \rightarrow x = x\] So the solution is all real numbers.

OpenStudy (ddcamp):

Because x=x will never be false.

OpenStudy (anonymous):

So Infinitely many solutions or no sollution

OpenStudy (ddcamp):

infinitely many